# Non-degenerate mixed maps and contact structures

### v1

@@ 2.1. Mixed maps
Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed map. We say that $F$ is non-degenerate with respect to the Newton boundaries $\Gamma(f^{1}),\dots,\Gamma(f^{k})$ if, for every vector $P$ of positive integers, the following condition is verified: at each point $p\in\mathbb{C}^{*n}$ such that $F_{P}(p)=0$, the differentials $Df_{P}^{1},{}\mkern 3.0mu\overline{\mkern-3.0muD}f_{P}^{1},\dots,Df_{P}^{k},{}\mkern 3.0mu\overline{\mkern-3.0muD}f_{P}^{k}$ do not satisfy a relation as in (1.2). In addition, we say that $F$ is strongly non-degenerate if the previous condition holds for any point $p\in\mathbb{C}^{*n}$.
@@
Recall the Siegel complete intersection map $\Psi_{\mathcal{F}}$ in Definition 1.7. Let $\Lambda=(\Lambda_{1},\dots,\Lambda_{n})$ the $n$-tuple of $\mathbb{C}^{k}$-vector defining this mixed ICIS. Let us suppose that for all $m$-tuples $(i_{1},\dots,i_{m})$ of integers in $\{1,\dots,n\}$, where $m\geq 2k$, the set $(\Lambda_{i_{1}},\dots,\Lambda_{i_{m}})$ is an admissible configuration. Then the mixed map $\Psi_{\mathcal{F}}$ becomes non-degenerate, since its face functions are the restrictions to subspaces $\mathbb{C}^{I}$ and $\Psi_{\mathcal{F}}^{I}=0$ has no solution in $\mathbb{C}^{*I}$ if $|I|<k$.
@@
We extend partial non-degeneracy for mixed maps. We remark that this definition for mixed functions on two variables was introduced in [3].
Definition 2.7.
Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed map. We say that $F$ is partially non-degenerate with respect to the Newton boundaries $\Gamma(f^{1}),\dots,\Gamma(f^{k})$ if, for every vector $P$ of positive integers and every nonempty subset $I\subset\{1,\dots,n\}$, the following condition is verified: at each point $p\in\mathbb{C}^{*n}$ such that $F_{P}^{I}(p)=0$, the differentials $(Df^{1})_{P}^{I}\;,\;({}\mkern 3.0mu\overline{\mkern-3.0muD}f^{1})_{P}^{I}\,,\dots,(Df^{k})_{P}^{I}\;,\;({}\mkern 3.0mu\overline{\mkern-3.0muD}f^{k})_{P}^{I}$ do not satisfy a relation as in (1.2).
Example 2.8.
Let $F:(\mathbb{C}^{3},0)\longrightarrow(\mathbb{C}^{2},0)$ be given by $F=(z_{1}+(z_{2}+z_{3})^{2},z_{1}^{2}+z_{2}^{2}+z_{3}^{2})$. The derivative $DF$ of $F$ consists of a constant and terms of order 1, so $DF=(DF)_{P}$ for every vector $P$ of positive integers. Moreover, for every non-empty $I\subsetneq\{1,2,3\}$, $(DF)^{I}_{P}$ does not have maximal rank only at points $p\in F_{P}^{-1}(0)$ in the complement of $\mathbb{C}^{*I}$. On the other hand, we may choose a vector $P$ so that the first coordinate function of $F_{P}$ is $(z_{2}+z_{3})^{2}$. In this case, $DF_{P}$ does not have maximal rank at points $p\in F_{P}^{-1}(0)\cap\mathbb{C}^{*3}$. Thus, $F$ is degenerate but partially non-degenerate.
@@
    If $F$ is non-degenarate, then $F^{I}$ is also non-degenerate, where $I\subset\{1,\dots,n\}$ is such that $f^{i,I}\not\equiv 0$ for every $i=1,\dots,k$.
@@
    If $F$ is (strongly, partially) non-degenerate and $\phi$ is a mixed covering, then the pullback $G=\phi^{*}F$ is also (strongly, partially) non-degenerate.
@@
    If $F$ is non-degenerate and convenient, then it is partially non-degenerate.
@@
For the first item, we follow the proof in [26, Proposition 7] for mixed functions. Let $P$ be a vector of positive integers and denote $\left(f^{i,I}\right)_{P}=f^{i,I}_{P}$ for every $i=1,\dots,k$. Let $Q_{i}=(q^{i}_{1},\dots,q^{i}_{n})$ be a vector such that $q^{i}_{j}=p_{i}$ if $i\in I$ and $q^{i}_{j}=v^{i}_{j}$ if $i\notin I$, where $v^{i}_{j}$ are positive integers. If $v^{i}_{j}$ are sufficiently large, then
$f^{i}_{Q_{i}}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=\left(f^{i,I}\right)_{P}(z_{I},{}\mkern 3.0mu\overline{\mkern-3.0muz}_{I}),$
@@
where $(z_{I},{}\mkern 3.0mu\overline{\mkern-3.0muz}_{I})=(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})\cap\mathbb{C}^{I}$. We may take $v^{i}=\max_{j}\{v^{i}_{j}\}$ and define $Q=(q_{1},\dots,q_{n})$ such that $q_{i}=p_{i}$ if $i\in I$ and $q_{i}=v_{i}$ if $i\notin I$. It follows that $F^{I}_{P}=F_{Q}$ and the non-degeneracy of $F$ is translated to the non-degeneracy of $F^{I}$.
In the second item, for each vector $P$ of positive integers one can see that $G_{P}=\phi^{*}F_{P}$. Since $\phi$ is a diffeomorphism on $\mathbb{C}^{*n}$, the assertion for (strong) non-degeneracy follows. On the other hand, for every non-empty subset $I\subset\{1,\dots,n\}$, the derivatives are related by $(DF)_{P}^{I}=(DG)_{P}^{I}\circ D\phi^{I}$. Since $\phi^{I}$ is a diffeomorphism of $\mathbb{C}^{*I}$ we conclude the proof for partially non-degenerate maps.
The last assertion is a consequence of the following remark (see [22, Proposition X1.7]). Let $f$ be a coordinate function of $F$. If $\left(f_{P}^{I}\right)_{z_{i}}\not\equiv 0$, then it is equal to $\left(f_{z_{i}}\right)^{I}_{P}$, where $I\subset\{1,\dots,n\}$ is non-empty. The analogous property holds for partial derivatives with respect to $\mkern 3.0mu\overline{\mkern-3.0muz}_{i}$. By the convenience hypothesis we can conclude the assertion by the first item.
@@
Item 2 of the proposition above is analogous to [28, Proposition 6] for mixed functions. One has that non-degeneracy property for holomorphic maps is a general condition by [16]. Therefore, this assertion provides several examples of non-degenerate mixed maps. The main result of this section relates non-degeneracy and ICIS properties.
@@
Let $F=(f^{1},\dots,f^{k}):\mathbb{C}^{n}\longrightarrow\mathbb{C}^{k}$ be a partially non-degenerate mixed map such that $V_{F}$ has positive dimension. Then $F$ is a mixed ICIS.
@@
On the contrary, by the Curve Selection Lemma and Proposition 1.1, there exists a real analytic curve $w:(0,1]\longrightarrow(\Sigma_{F}\cap V_{F})\setminus\{0\}$ such that $\lim_{t\to 0}w(t)=0$, and real analytic curves $\alpha_{1},\dots,\alpha_{k}\subset\mathbb{C}$ non-simultaneously vanishing such that
$\alpha_{1}(t){}\mkern 3.0mu\overline{\mkern-3.0muD}f^{1}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right)+\alpha_{k}(t){}\mkern 3.0mu\overline{\mkern-3.0muD}f^{k}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right)={}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{1}(t){}\mkern 3.0mu\overline{\mkern-3.0muDf^{1}}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right)+\dots+{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{k}(t){}\mkern 3.0mu\overline{\mkern-3.0muDf^{k}}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right).$
Let us suppose that $I=\{i:w_{i}(t)\not\equiv 0\}$ and without loss of generality, $\alpha_{i}(t)\not\equiv 0$ for all $i=1,\dots,k$. Consider the real analytic expansion of the curves:
@@
$\alpha_{i}(t)=c_{i}t^{q_{i}}+o(t),\;c_{i}\in\mathbb{C}^{*},q_{i}\geq 1,$
for all $i\in I$, where $o(t)$ denotes higher order degree terms. Let $d_{i}$ be the weighted homogeneous degree of the face functions $f^{i}_{P}$, for $i=1,\dots,k$. For each $i$, the following equations hold:

### v3

@@ 2.1. Mixed maps
Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed map germ. We say that $F$ is non-degenerate with respect to the Newton boundaries $\Gamma(f^{1}),\dots,\Gamma(f^{k})$ if, for every vector $P$ of positive integers, the following condition is verified: at each point $p\in\mathbb{C}^{*n}$ such that $F_{P}(p)=0$, the differentials $\mkern 3.0mu\overline{\mkern-3.0muDf_{P}^{1}},{}\mkern 3.0mu\overline{\mkern-3.0muD}f_{P}^{1},\dots,{}\mkern 3.0mu\overline{\mkern-3.0muDf_{P}^{k}},{}\mkern 3.0mu\overline{\mkern-3.0muD}f_{P}^{k}$ do not satisfy a relation as in (1.2). In addition, we say that $F$ is strongly non-degenerate if the previous condition holds for any point $p\in\mathbb{C}^{*n}$.
@@
Recall the Siegel complete intersection map $\Psi_{\mathcal{F}}$ in Definition 1.7. Let $\Lambda=(\Lambda_{1},\dots,\Lambda_{n})$ be the $n$-tuple of $\mathbb{C}^{k}$-vector defining this mixed ICIS. Let us suppose that for all $m$-tuples $(i_{1},\dots,i_{m})$ of integers in $\{1,\dots,n\}$, where $m>2k$, the set $(\Lambda_{i_{1}},\dots,\Lambda_{i_{m}})$ is an admissible configuration. Then the mixed map $\Psi_{\mathcal{F}}$ becomes non-degenerate for every face $\Delta$ such that $F_{\Delta}$ is a map on $m>2k$ variables, since its face functions are the restrictions to subspaces $\mathbb{C}^{I}$.
@@
Remark 2.8.
Let us consider $n\leq k$ in Definition 2.5. In the case $n<k$, we have that $\mathbb{C}^{n}\subset\Sigma_{F_{P}}$ for every vector $P$ of positive integers, and thus $F_{P}^{-1}(0)\cap\mathbb{C}^{*n}=\emptyset$. If $n=k$, and there exists $p\in F_{P}^{-1}(0)\cap\mathbb{C}^{*n}$, then $F_{P}$ is a diffeomorphism in a small neighborhood of $p$. Since $F_{P}(0)=0$, we obtain a contradiction. We conclude that $F_{P}^{-1}(0)\cap\mathbb{C}^{*n}=\emptyset$ for every $P$. Since the applications of this definition refer to points in these intersections, we exclude this case. In particular, when we consider the non-degeneracy for the restrictions $F^{I}:(\mathbb{C}^{I},0)\longrightarrow(\mathbb{C}^{k},0)$ of $F$ to subspaces $\mathbb{C}^{I}$, we suppose $|I|>k$.
We prove some properties of non-degenerate maps. The third item of the following result illustrates how the mixed coverings in Subsection 1.3 play a special role in the construction of mixed maps with particular properties. Let us fix some notations now. Let $I\subset\{1,\dots,n\}$ such that $|I|>k$ and $J_{I}=\{j:f^{j,I}\not\equiv 0\}$. We denote by $F^{I}_{J}:(\mathbb{C}^{I},0)\longrightarrow(\mathbb{C}^{J},0)$ the mixed map germ whose coordinate functions are those $f^{j}$ for which $f^{j,I}\not\equiv 0$.
@@
    If $F$ is (strongly) non-degenarate, then $F^{I}$ is also (strongly) non-degenerate, where $I\subset\{1,\dots,n\}$ is such that $|I|>k$ and $f^{i,I}\not\equiv 0$ for every $i=1,\dots,k$.
@@
    If $F$ is strongly non-degenerate, then $F^{I}_{J}$ is also strongly non-degenerate for every $I\subset\{1,\dots,n\}$ such that $|I|>k$ and $F^{I}\not\equiv 0$.
@@
    If $F$ is (strongly) non-degenerate and $\phi$ is a mixed covering, then the pullback $G=\phi^{*}F$ is also (strongly) non-degenerate.
@@
For the first item, we follow the proof in [30, Proposition 7] for mixed functions. Let $P$ be a vector of positive integers and denote $\left(f^{j,I}\right)_{P}=f^{j,I}_{P}$ for every $j=1,\dots,k$. Let $Q_{j}=(q^{j}_{1},\dots,q^{j}_{n})$ be a vector such that $q^{j}_{i}=p_{i}$ if $i\in I$ and $q^{j}_{i}=v_{i}^{j}$ if $i\notin I$, where $v^{j}_{i}$ is a positive integer. If $v^{j}_{i}$ is sufficiently large for every $i\notin I$, then
$f^{j}_{Q_{j}}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=\left(f^{j,I}\right)_{P}(z_{I},{}\mkern 3.0mu\overline{\mkern-3.0muz}_{I}),$
@@
where $(z_{I},{}\mkern 3.0mu\overline{\mkern-3.0muz}_{I})=(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})\cap\mathbb{C}^{I}$. We may take $v_{i}=\max_{j}\{v^{j}_{i}\}$ and define $Q=(q_{1},\dots,q_{n})$ such that $q_{i}=p_{i}$ if $i\in I$ and $q_{i}=v_{i}$ if $i\notin I$. It follows that $F^{I}_{P}=F_{Q}$ and the non-degeneracy of $F$ is translated to the non-degeneracy of $F^{I}$.
Consider the second claim. Let $J=\{j_{1},\dots,j_{m}\}$ be the indices such that $f^{j_{1},I},\dots,f^{j_{m},I}\not\equiv 0$. By the hypothesis $F^{I}\not\equiv 0$, the set $J$ is non-empty. Define $F^{I}_{J}=(f^{j_{1},I},\dots,f^{j_{m},I}):(\mathbb{C}^{I},0)\longrightarrow(\mathbb{C}^{J},0)$. We shall prove that $F^{I}_{J}$ is strongly non-degenerate. Let $P$ be a vector of positive integers such that $p\in\mathbb{C}^{*I}$ is a critical point of $\left(F^{I}_{J}\right)_{P}$. This implies that the gradients $\mkern 3.0mu\overline{\mkern-3.0muD}\left(f^{j_{r},I}\right)_{P}$, $\mkern 3.0mu\overline{\mkern-3.0muD\left(f^{j_{r},I}\right)}_{P}$, where $r=1,\dots,m$, are linearly dependent at $p$. By the previous item, there exists a vector $Q$ such that $f^{j_{r},I}_{P}=f^{j_{r}}_{Q}$ for every $r$. Thus, $q=(p,\tilde{p})\in\mathbb{C}^{*n}$ becomes a critical point of $F_{Q}$ for all $\tilde{p}\in\mathbb{C}^{*I^{C}}$, which is a contradiction.
In the third item, for each vector $P$ of positive integers, one can see that $G_{\tilde{P}}=\phi^{*}F_{P}$, where $\tilde{P}\in\mathbb{Q}^{n}$ is a vector whose coordinates $\tilde{p}_{i}$ are normalized by the exponents $a_{i},b_{i}$ of the mixed covering $\phi=\phi_{\textbf{a},\textbf{b}}$ (see [32, Proposition 6]). Since $\phi$ is a diffeomorphism on a sufficiently small neighborhood of $\mathbb{C}^{*n}$, the assertion follows. ∎
@@
Items 1 and 2 of the above proposition are analogous to [32, Proposition 6] for mixed functions. One has that the non-degeneracy property for holomorphic maps is a general condition by [19]. Therefore, this assertion provides several examples of non-degenerate mixed maps.
In the applications of the Newton non-degeneracy property, one needs to a proper notion of convenience, which is related with the real analytic curves provided by the Curve Selection Lemma. An equivalent description of a convenient mixed function germ $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ is the following. For each non-empty subset $I\subset\{1,\dots,n\}$, the restriction $f^{I}\not\equiv 0$. In the case of mixed maps, we will be able to avoid this condition strictly for each coordinate function. In the sequel, we motivate our conditions.
In virtue of Proposition 2.9, Remark 2.8 implies the following. Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a non-degenerate mixed map germ such that $f^{j}$ is convenient for every $j=1,\dots,k$. Let $I\subset\{1,\dots,n\}$ be a non-empty subset such that $|I|\leq k$ and $P$ a vector of positive integers. Let $Q$ be a second vector of positive integers such that $F^{I}_{P}=F_{Q}$. As we argued in Remark 2.8, one obtains that $V_{F_{Q}}\cap\mathbb{C}^{*I}=\emptyset$. This is of particular interest, since the main result of this section concerns the singular points of mixed varieties. The next proposition motivates our first definition of convenience.
Proposition 2.10.
Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed map germ. Let $I,J\subset\{1,\dots,n\}$ be non-empty subsets. Then the following properties hold.
  1. (1)
    The condition $V_{F}\cap\mathbb{C}^{I}=\{0\}$ for every $I$ such that $|I|=k$ is equivalent to $V_{F}\cap\mathbb{C}^{*J}=\emptyset$ for every $J$ with $|J|\leq k$.
  2. (2)
    If $F$ is non-degenerate and $f^{j}$ is convenient for each $j$, then $V_{F}\cap\mathbb{C}^{I}=\{0\}$ for every $I$ such that $|I|=k$.
Proof.
The first assertion follows from the fact that every $J$ with $|J|\leq k$ is a subset of some $I$ with $|I|=k$. For the second item, if the statement is not true for some $I$ with $|I|=k$, we apply the Curve Selection Lemma to obtain a real analytic curve $w:(0,1]\longrightarrow\mathbb{C}^{n}$ such that $\lim_{t\to 0}w(t)=0$ and $w(t)\subset V_{F}\cap\mathbb{C}^{I}$. Let $J=\{j:w_{j}(t)\not\equiv 0\}\subset I$ and $w_{j}=a_{j}t^{p_{j}}+o(t)$ the analytic expansion of each coordinate $w_{j}$, where $j\in J$, $a_{j}\neq 0$, and $o(t)$ denotes higher order terms. Since $f^{j,J}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))\equiv 0$, we obtain that the face functions $(f^{j,J})_{P}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})=0$ for every $j$, where the coordinates of $P$ are $p_{j}$ for $j\in J$ (see the proof of Theorem 2.13 for details). Since $f^{j}$ is convenient, by the first item of Proposition 2.9, there exists a vector $Q$ such that $F^{J}_{P}=F_{Q}$. We conclude that $a\in V_{F_{Q}}\cap\mathbb{C}^{*J}$, which contradicts the discussion in the previous paragraph. ∎
Example 2.11.
Consider the map germ $F=(f^{1},f^{2}):(\mathbb{C}^{4},0)\longrightarrow(\mathbb{C}^{2},0)$ given by:
$F(z_{1},z_{2},z_{3},z_{4})=(z_{1}^{2}+z_{2}^{2}+z_{3}^{2}+z_{3}z_{4}+T(z,{}\mkern 3.0mu\overline{\mkern-3.0muz}),-z_{1}^{2}+z_{2}^{2}-2z_{3}^{2}+z_{4}^{2}),$
where $T(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ is a mixed function consisting of monomials above the Newton polyhedron of $z_{1}^{2}+z_{2}^{2}+z_{3}^{2}+z_{3}z_{4}$. Notice that $f^{1}$ is not convenient, but $f^{1,I}\not\equiv 0$ for all $I\subset\{1,2,3,4\}$ such that $|I|>2$. We also observe that the compact faces of $f^{1}$ do not depend on the terms of $T(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$. Moreover, a direct computation shows that $V_{F}\cap\mathbb{C}^{I}=0$ if $|I|=2$, and that $F$ is also non-degenerate. We shall see that these conditions ensure the ICIS property.
Definition 2.12.
Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed map germ. We say that $F$ is convenient if the following conditions hold for every $j=1,\dots,k$ and non-empty $I\subset\{1,\dots,n\}$:
  1. (1)
    If $|I|>k$, then $f^{j,I}\not\equiv 0$.
  2. (2)
    If $|I|=k$, then $V_{F}\cap\mathbb{C}^{I}=\{0\}$.
In other words, convenient mixed maps are those for which the zero set does not contain small coordinate subspaces of $\mathbb{C}^{n}$. From the second item of Proposition 2.10, one has that mixed map germs whose all coordinate functions are convenient, are convenient as well. As the Curve Selection Lemma suggests, these subspaces are related to the loss of the non-degenerate property in the restrictions $F^{I}$ of $F$. In the next subsection, we study in detail a non-degenerate mixed map germ whose all coordinate functions are convenient. A number of examples can be constructed through mixed coverings, as we shall remark in Proposition 2.18. We can state now the main result of this section.
@@
Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a non-degenerate and convenient mixed map germ such that $V_{F}$ has positive dimension. Then $F$ is a mixed ICIS.
@@
We follow a standard argument of the literature (see, for instance, [28, Section 8.3]). If the statement is false, then by the Curve Selection Lemma and Proposition 1.1, there exist real analytic curves $w(t)\subset V_{F}\cap\Sigma_{F}$, $\alpha_{1}(t),\dots,\alpha_{k}(t)\subset\mathbb{C}$ non-simultaneously vanishing such that:
$\alpha_{1}(t){}\mkern 3.0mu\overline{\mkern-3.0muD}f^{1}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right)+\dots+\alpha_{k}(t){}\mkern 3.0mu\overline{\mkern-3.0muD}f^{k}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right)=-{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{1}(t){}\mkern 3.0mu\overline{\mkern-3.0muDf^{1}}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right)-\dots-{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{k}(t){}\mkern 3.0mu\overline{\mkern-3.0muDf^{k}}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right).$
Let $I=\{i:w_{i}(t)\not\equiv 0\}$ and notice that $|I|>k$ by the convenience. Consider the real analytic expansion of the curves:
@@
$\alpha_{j}(t)=c_{j}t^{q_{j}}+o(t),\;c_{j}\in\mathbb{C}^{*},q_{j}\geq 1,$
for all $i\in I$ and $j$ such that $\alpha_{j}\not\equiv 0$, where $o(t)$ denotes higher order degree terms. We obtain that:
(2.1) $\begin{split}\frac{\partial f^{j,I}}{\partial z_{i}}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))&=\frac{\partial f_{P}^{j,I}}{\partial z_{i}}t^{d_{j}-p_{i}}+o(t),\\ \frac{\partial f^{j,I}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))&=\frac{\partial f_{P}^{j,I}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}t^{d_{j}-p_{i}}+o(t),\end{split}$
where $d_{j}$ is the weighted homogeneous degree of the face function $f^{j,I}_{P}$ of $f^{j,I}$. Replacing these expansions in the relation above, for each $i$, we have: